OpenStudy (nuccioreggie):

Solve x^3 = 64 over 27

OpenStudy (nuccioreggie):

@LegendarySadist

OpenStudy (nuccioreggie):

@Nnesha

OpenStudy (anonymous):

\[x^3=\frac{ 64 }{ 27 },27x^3-64=0\] \[\left( 3x \right)^3-3^3=0\] \[a^3-b^3=\left( a-b \right)\left( a^2+ab+b^2 \right)\]

OpenStudy (nuccioreggie):

So is it. C

OpenStudy (anonymous):

i don;t see options.

OpenStudy (nuccioreggie):

Solve x3 = 64 over 27. ±8 over 3 8 over 3 ±4 over 3 4 over 3

OpenStudy (anonymous):

correction \[\left( 3x \right)^3-4^3=0\]only real value is 3x-4=0 \[x=\frac{ 4 }{ 3 }\]

OpenStudy (nuccioreggie):

So its D right

OpenStudy (loser66):

\(64 = ?^3\)

OpenStudy (loser66):

\(27=??^3\)

OpenStudy (loser66):

\(x^3 =\dfrac{64}{27}=\dfrac{?^3}{??^3}=(\dfrac{?}{??})^3\) hence \(x= \dfrac{?}{??}\) your duty is find out what are ? and ??

OpenStudy (nuccioreggie):

@Loser66. D

OpenStudy (nuccioreggie):

So its D right @Nnesha